OpenStudy (anonymous):

Solve the following system of equations. 2x – 3y + z = –1 3x + 2y + 2z = –1 x – y – 3z = –4

OpenStudy (anonymous):

if you use elimination 6x-2y=-6 solve for a variable subtract 6x -2y=-6x-6 divide by -2 y=3x+3 substitute into the second equation 3x+2y+2z=-1 3x+2(3x+3)+2z=-1 2x+6x+6+2z=-1 combine like terms 8x+6+2z=-1 subtract 6 8x+2z=-7 subtract 8x 2z=8x+7 divide by 2 z=4x+(7/2) substitute the y and z solutions into the third equation x – y – 3z = –4 x-(3x+3)-3(4x+(7/2)=-4 distribute x-3x-3-12x-(21/2)=-4 combine like terms -14x-(27/2)=-4 add 27/2 -14x=(19/2) divide by -14 x=(-19/28) substitute x in for the y and z equations y=3(-19/28)+3 PEMDAS y=27/28 z=4(-19/28)+(7/2) PEMDAS z=11/14 so your coordinate triple is \[\left( \frac{ -19 }{ 28 },\frac{ 27 }{ 28 },\frac{ 11 }{ 14 } \right)\]

OpenStudy (anonymous):

sorry it's quite long there is a lot of work for this kind of thing